OpenStudy (anonymous):

Find F^-1 of the equation F(x)=2x+3

OpenStudy (anonymous):

To find the inverse, as you want to, swap x and y. \[y = 2x+3 \] \[x = 2y+3 \] \[x-3 = 2y\] \[y = \frac x2 - \frac 32\]

OpenStudy (anonymous):

ok awesome thank you!